.If 12 years are added to 2/3 age of Rani, she will be three years older than today. What is Rani’s present age?
Answer by hovuquocan1997(83)

Let set Rani's age = x and rewrite the question
If 12 are added to 2/3 of x, the final result will be x+3
So we have the equation:
12 + 2/3x = x + 3
Now we just solve the equation :D
12 + 2/3x = x + 3 -> multiply 3 to both sides to get rid of the bottom 3
36 + 2x = 3x + 9
27 = x
So Rani's present age is 27 years old.
:D